A mother's age is now 3 times that of her son. 8 years ago, the mother's age was greater than 4 times that of her son by 4. Find the son's present age.
Answer by ankor@dixie-net.com(22740)

let m = Mom's age
let s = son's age:
write an equation for each statement
:
A mother's age is now 3 times that of her son.
m = 3s
:
8 years ago, the mother's age was greater than 4 times that of her son by 4.
m - 8 = 4(s - 8) + 4
m - 8 = 4s - 32 + 4
m = 4s - 28 + 8
m = 4s - 20
Replace m with 3s from the 1st statement
3s = 4s - 20
20 = 4s - 3s
20 = s
:
Find the son's present age. 20 yrs
:
:
:
Confirm this. Mom is 3(20) = 60 yrs
Check this in the 2nd statement
"8 years ago, the mother's age was greater than 4 times that of her son by 4."
60 - 8 = 4(20 - 8) + 4
52 = 4(12) + 4
52 = 48 + 4